### v1

Abstract.
We introduce and investigate a class of ring ideals, termed ring $\operatorname{M}$-ideals, inspired by the Alfsen–Effros theory of $\operatorname{M}$-ideals in Banach spaces. We show that $\operatorname{M}$-ideals extend the classical notion of essential ideals and subsume them as a subclass. The central theorem provides a full characterization: an ideal is an $\operatorname{M}$-ideal if and only if it is either essential or relatively irreducible. This dichotomy reveals the abundant and diverse nature of $\operatorname{M}$-ideals, encompassing both essential and minimal ideals, and admits natural generalizations in rings beyond the commutative and unital settings.
We systematically study the algebraic stability of $\operatorname{M}$-ideals under standard constructions such as intersection, quotient, direct product, and Morita equivalence and establish their behavior in topological rings and operator algebras. In certain rings such as $\mathds{Z}_{n}$, and C*-algebras, we completely classify $\operatorname{M}$-ideals and relate them to algebraically minimal projections and central idempotents. The ring $\operatorname{M}$-ideals in $C(K)$ are shown to be precisely the essential ideals or those minimal ideals corresponding to isolated points.
Structurally, we show that the absence of proper $\operatorname{M}$-ideals characterizes simplicity, while rings in which every proper $\operatorname{M}$-ideal is a direct summand must decompose as finite direct sums of simple rings. In closing, we introduce the notion of $\operatorname{M}$-complements, drawing an analogy with essential extensions in module theory, and demonstrate their existence.
Key words and phrases:
$\operatorname{M}$-ideal, essential ideal, simple ring, minimal ideal, socle, Morita equivalence, ring decomposition
1991 Mathematics Subject Classification:
16D25, 16D70, 16D90, 46L05, 46B28
1. Introduction
The $\operatorname{M}$-ideals in a Banach space are one of the most important tools in the isometric theory of Banach spaces and constitute a formidable technology [9]. Drawing an analogy from their beautiful characterization by Alfsen and Effros, we are led to define a class of ideals in rings. It turns out that this class contains the essential ideals. Recall that an ideal $J$ of a ring $R$ is called essential if $J\cap I\neq 0$ for all nonzero ideals $I$ of $R$. As a generalization thereof, essential submodules play a pivotal role in the theory of injective modules via the notion of injective hulls. Borrowing ideas from various concepts related to essential modules, we introduce an analogous notion in the context of $\operatorname{M}$-ideals.
The aim of this paper is twofold: on the one hand, to systematically develop a theory of ring $\operatorname{M}$-ideals as a generalization of essential ideals; on the other hand, to relate this class of ideals to the corresponding notion in Banach spaces and C*-algebras. The subtitle of our paper, in some sense, should be “from Banach spaces back to rings" because part of Alfsen and Effros’ motivation was to ‘generalize’ to Banach spaces the usual ‘calculus’ of ideals in rings [10, 12]. By now, however, $\operatorname{M}$-ideals in Banach spaces have a beautiful and formidable theory of their own [9], and some of these ideas may in turn be applied in algebra.
We now describe the structure of our paper. We investigate the stability of this class under the usual constructions with ring ideals, such as intersections, quotients, etc. We establish that $\operatorname{M}$-ideals strictly generalize essential ideals (Proposition 2.3), and provide a complete characterization: an ideal is an $\operatorname{M}$-ideal if and only if it is essential or relatively irreducible (Theorem 3.1). This characterization allows us to classify $\operatorname{M}$-ideals explicitly in important classes of rings, such as $\mathds{Z}_{n}$ (Theorem 3.4), $C(K)$ (Theorem 3.7), and unital C*-algebras. Indeed, the nonzero ring $\operatorname{M}$-ideals in a C*-algebra are precisely either the essential ideals or the ideals whose closure is a prime C*-algebra (See Corollary 4.3. Prime C*-algebras are well understood, and have a large theory, e.g. [3].) We show that $\operatorname{M}$-ideals are not preserved under direct sums or tensor products. We show that for any ring with proper ideals necessarily has proper $\operatorname{M}$-ideals (Corollary 5.4). Further, we show that Morita equivalence preserves M-ideals (Theorem 5.5) and that every ideal is contained in a maximal $\operatorname{M}$-ideal unless it is a direct summand. From a structural standpoint, we prove that the absence of proper $\operatorname{M}$-ideals characterizes simple rings and that rings in which every proper $\operatorname{M}$-ideal is a direct summand must decompose as finite direct sums of simple rings (Theorem 5.3). Motivated by analogies with essential extensions in module theory, we introduce the notion of $\operatorname{M}$-complements and establish their universal existence (Theorem 5.9), thereby enriching the interplay between ideal theory and ring decomposition.
2. Basic Properties
According to Theorem 5.8 in [2], an ideal $J$ of a Banach space is an $\operatorname{M}$-ideal if and only if the following condition holds: if $B_{1}$, $\ldots$, $B_{n}$ are open balls with $B_{1}\cap\cdots\cap B_{n}\neq\emptyset$ and $B_{i}\cap J\neq\emptyset$ for all $i$, then $B_{1}\cap\cdots\cap B_{n}\cap J\neq\emptyset$. Motivated by this characterization, we propose the following definition for an $\operatorname{M}$-ideal of a ring.
Definition 2.1.
[…]
$J\cap\left(\bigcap_{k=1}^{n}I_{k}\right)\neq 0.$
For brevity, henceforth we shall usually simply use the expression $\operatorname{M}$-ideal for ‘ring $\operatorname{M}$-ideal’, and only add the ‘ring’ adjective where there is a possibility of confusion with the Banach space notion (that is, in Banach space settings).
All of our rings will be supposed to have identity for simplicity, although this is often not needed, and Definition 2.1 makes perfect sense for nonunital rings. However, in Corollary 2.11, we show that a subset $J$ of a nonunital ring $R$ is an $\operatorname{M}$-ideal if and only if $J$ is an $\operatorname{M}$-ideal in the unitization of $R$. So, we may safely avoid nonunital rings. In a few of the results, we will require the ring to be commutative.
Although our choice of taking $n$ ideals $I_{1}$, $\ldots$, $I_{n}$ in Definition 2.1 was motivated by the characterization of $\operatorname{M}$-ideals of Banach spaces, it is, however, possible to restrict it to $n=2$.
Proposition 2.1.
In any ring, the following are equivalent:
  1. (1)
    $J$ is an $\operatorname{M}$-ideal.
  2. (2)
    For every nonzero ideal $I_{1},$ $I_{2}$ of $R$ such that $I_{1}\cap I_{2}\neq 0$, the conditions $J\cap I_{1}\neq 0$ and $J\cap I_{2}\neq 0$ imply that $J\cap(I_{1}\cap I_{2})\neq 0$.
Proof.
Obviously, (1) implies (2). To show (2) implies (1), we use induction on $n$. By (2), the property holds for $n=2$. Assume that the property holds for $n$. Let $I_{1},\ldots,I_{n+1}$ be a collection of nonzero ideals of $R$ such that $\bigcap_{k=1}^{n+1}I_{k}\neq 0$, and for each $k$ we have $J\cap I_{k}\neq 0$. By inductive hypothesis, we have $J\cap(\bigcap_{k=1}^{n}I_{k})\neq 0$. Note that
$\left(\bigcap_{k=1}^{n}I_{k}\right)\cap I_{n+1}=\bigcap_{k=1}^{n+1}I_{k}\neq 0$
and $J\cap I_{n+1}\neq 0$. So by assumption, we conclude that
$J\cap\left(\bigcap_{k=1}^{n+1}I_{k}\right)=J\cap\left(\bigcap_{k=1}^{n}I_{k}\right)\cap I_{n+1}\neq 0.\qed$
Remark 2.2.
According to [2, Theorem 5.9], the characterization of $\operatorname{M}$-ideals of Banach spaces can be restricted to three open balls, but not to two. We just saw that the situation is different for rings.
We note that if the ring $R$ is a Banach space, then an $\operatorname{M}$-ideal of $R$ in the sense above need not be an $\operatorname{M}$-ideal of $R$ in the Banach space sense, nor vice versa. Indeed, suppose that $R=C(K)$, where an $\operatorname{M}$-ideal $J$ in the latter sense must be the subspace $J_{E}$ of functions vanishing on a closed subset $E\subset K$ [9]. Similarly, take $I_{k}=J_{E_{k}}$ for a closed subsets $E_{1},E_{2}$ in $K$. Note that $J_{\bigcup_{k}\,E_{k}}=\bigcap_{k}\,J_{k}$. It is easy to find examples where $\bigcup_{k}\,E_{k}\neq K$, so that $\bigcap_{k=1}^{2}I_{k}\neq 0$, and $E\cup E_{k}\neq K$ for each $k$, but $E\cup(\bigcup_{k}E_{k})=K$. That is,
$J\cap\left(\bigcap_{k=1}^{2}I_{k}\right)=0.$
indeed, this happens in $K=[0,1]$. Thus $J$ is not a ring $M$-ideal. See also Proposition 3.7.
Conversely, consider $A=l^{\infty}$, with $J$ the finitely supported sequences. This intersects every nonzero ideal (the noncommutative analogue is contained in every nonzero ideal), so it is a ring $\operatorname{M}$-ideal of $R$ but not a Banach space $\operatorname{M}$-ideal since it is not closed.
As indicated in the introduction, our notion of $\operatorname{M}$-ideals are a generalization of essential ideals; hence, naturally, these are our first examples.
Proposition 2.3.
Every essential ideal of a ring is an $\operatorname{M}$-ideal.
Proof.
Suppose that $J$ is an essential ideal (and hence nonzero) of ring $R$. Let $I_{1}$ and $I_{2}$ be two nonzero ideals of $R$ such that $I_{1}\cap I_{2}\neq 0$ and $J\cap I_{k}\neq 0$ for each $k$. Since $J$ is essential, we have $J\cap(I_{1}\cap I_{2})\neq 0$. ∎
Remark 2.4.
The reader may wonder whether every $\operatorname{M}$-ideal is essential. To see this is not true, consider the ring $\mathds{Z}_{12}$ and the ideal $J:=(3)$. Take $I_{1}:=(2)$ and $I_{2}:=(6)$. It is easy to see that $(3)$ is an $\operatorname{M}$-ideal. Note that we cannot include $(4)$ as $I_{3}$. However, $(3)$ is not an essential ideal because $(3)\cap(4)=(0).$
We claim that $\operatorname{M}$-ideals are abundant, and the following proposition will validate this. Of course, we shall see more examples as we proceed.
Proposition 2.5.
In a ring $R$, the following hold.
  1. (1)
    The zero ideal is an $\operatorname{M}$-ideal.
  2. (2)
    The maximal ideal of a nontrivial local ring is an $\operatorname{M}$-ideal.
  3. (3)
    Every ideal of an integral domain is an $\operatorname{M}$-ideal.
  4. (4)
    If $R$ is an algebra and $e$ is a central element which is an ‘algebraically minimal idempotent’ in $R$ (so $eR=\operatorname{{\mathds{F}}}e$) then $Re$ is an $\operatorname{M}$-ideal.
  5. (5)
    If $R$ is a ring and $e$ is a central element which satisfies $eR=R^{-1}e$ (or $eR=({\mathcal{Z}}(R))^{-1}e$) then $Re$ is an $\operatorname{M}$-ideal.
  6. (6)
    If $J$ is a minimal ideal in $R$, then $J$ is an $\operatorname{M}$-ideal.
Proof.
The assertion (1) is vacuously true. The properties (2) and (3) follow from Proposition 2.3. For (4), observe that if $R$ is an algebra and $e$ is a central element which is a minimal idempotent in $R$, and $J$ is an ideal in $R$, then $J\cap eR\neq 0$ if and only if $e\in J$. Now it is clear $eR$ is an $\operatorname{M}$-ideal. The proof of (5) is similar to the previous proof. An ideal $I$ intersects $eR$ if and only if it contains $er$ for an invertible $r$, and so if and only if $e\in I$. Now it is clear $eR$ is an $\operatorname{M}$-ideal. The property (6) is in fact generalizes the last couple of items. If $J$ is a simple ideal in $R$ and $K$ is another nonzero ideal, then $J\cap K\neq 0$ implies that $J\cap K=J$. The result is clear from this. ∎
Remark 2.6.
Note that a zero ideal of a ring, by definition, is not essential, but it is an $\operatorname{M}$-ideal. In the ring $\mathds{Z}$ of integers, for each prime $p$, the ideal $(p)$ is an essential ideal, and hence, an $\operatorname{M}$-ideal by Proposition 2.5(3). However, $\bigcap_{p}(p)=(0)$ is not essential but is an $\operatorname{M}$-ideal. Although $\operatorname{M}$-ideals are a generalization of essential ideals, interestingly enough, minimal ideals are very far from being essential.
Jacobson radicals are generally not essential. For example, let us consider the ring $\mathds{Z}_{24}$. In this ring, $\mathrm{Jac}(\mathds{Z}_{24})=(6)$, an $\operatorname{M}$-ideal; however, it is not essential because $(6)\cap(8)=(0)$.
Proposition 2.7.
The intersection of a finite number of $\operatorname{M}$-ideals is an $\operatorname{M}$-ideal.
Proof.
[…]
$\bigcap_{k=1}^{2}(J_{2}\cap I_{k})=J_{2}\cap(I_{1}\cap I_{2})\neq 0,$
and $J_{1}\cap(J_{2}\cap I_{k})\neq 0$ for each $k$. Since $J_{1}$ is also an $\operatorname{M}$-ideal, we have that
$(J_{1}\cap J_{2})\cap(I_{1}\cap I_{2})=J_{1}\cap\left(\bigcap_{k=1}^{2}(J_{2}\cap I_{k})\right)\neq 0.\qed$
The intersection of an infinite number of $\operatorname{M}$-ideals is not necessarily an $\operatorname{M}$-ideal. For example, consider the intersection of the (essential) ideals $J_{r}$ in $C([0,1])$ of functions vanishing at a rational point $r$ in $[0,1]$.
Proposition 2.8.
If $J$ is an $\operatorname{M}$-ideal of $R$, and $K$ is an ideal of $R$ which is contained in $J$, then $J/K$ is an $\operatorname{M}$-ideal of $R/K$.
Proof.
Suppose that $I_{1},I_{2}$ are ideals in $R/K$ which intersect each other and intersect $J/K$. One easily checks that $q_{K}^{-1}(I_{1}),q_{K}^{-1}(I_{2})$ are ideals in $R$ which intersect each other and intersect $J$. Thus $J\cap q_{K}^{-1}(I_{1})\cap q_{K}^{-1}(I_{2})\neq 0$. If $x$ is in the latter intersection, then $q_{K}(x)\in(J/K)\cap I_{1}\cap I_{2}$. ∎
Proposition 2.9.
The direct product of a finite number of $\operatorname{M}$-ideals is an $\operatorname{M}$-ideal. An $\operatorname{M}$-ideal of a direct product of a finite number of rings is a direct product of $\operatorname{M}$-ideals.
Proof.
This relies on the well-known fact that the ideals of a direct product of a finite number of rings are exactly the direct products of ideals. We leave the first assertion as an exercise using this fact.
Suppose that $I$ is an $\operatorname{M}$-ideal of $\prod_{i=1}^{n}\,R_{k}$. Then by the fact above, $I=\prod_{i=1}^{n}\,I_{k}$. Suppose that $J_{1}$ and $K_{1}$ are ideals in $R_{1}$ with $J_{1}\cap K_{1},J_{1}\cap I_{1},I_{1}\cap K_{1}$ nonzero. Then the intersection of any two of
$J_{1}\times\left(\prod_{i=2}^{n}\,I_{k}\right),\;\;\;K_{1}\times\left(\prod_{i=2}^{n}\,I_{k}\right),\;\;\;I=I_{1}\times\left(\prod_{i=2}^{n}\,I_{k}\right)$
is nonzero. Thus, the intersection of these three is nonzero. Thus $I_{1}\cap J_{1}\cap K_{1}\neq 0$. Hence $I_{1}$ is an $\operatorname{M}$-ideal, and similarly so is $I_{2},\cdots,I_{n}$. Thus $I$ is a direct product of $\operatorname{M}$-ideals. ∎
Remark 2.10.
Following on from the last result, one may ask if an $\operatorname{M}$-ideal of a subdirect product $R$ of a finite number of rings $R_{i}$ is a subdirect product of $\operatorname{M}$-ideals of $R_{i}$? However, this is false. If $n=p_{1}\cdots p_{m}$ is a product of primes, then $\mathds{Z}_{n}$ is a subdirect product of the simple rings $\mathds{Z}_{p_{i}}$. However, it is easy to see that the $\operatorname{M}$-ideal $(2)$ of $\mathds{Z}_{6}$ is not a subdirect product of $\mathds{Z}_{2}$ and $\mathds{Z}_{3}$.
In addition, it is also false that in the setting of the last paragraph of a subdirect product $R$ of rings $R_{i}$, $\operatorname{M}$-ideals $I_{k}$ of $R_{k}$ correspond to an $\operatorname{M}$-ideal of $R$. To see this consider $\mathds{Z}_{90}$ as the subdirect product of $\mathds{Z}_{2}$, $\mathds{Z}_{3}$ and $\mathds{Z}_{5}$, as in the last paragraph. Then $(2)$ is a subdirect product of the $\operatorname{M}$-ideals $(2),(0),(0)$ of the latter three rings. However $(2)$ is not an $\operatorname{M}$-ideal of $\mathds{Z}_{90}$, as may be seen for example from Theorem 3.4.
If we allow nonunital rings, several other items in the ‘calculus of Banach space $\operatorname{M}$-ideals’ carry over. For example, suppose that $J$ is an ideal in $R$, and $K$ is an ideal of $J$. Then if $K$ is an $\operatorname{M}$-ideal of $R$, then $K$ is clearly an $\operatorname{M}$-ideal in $J$. Conversely, if $J$ is a ring $\operatorname{M}$-ideal of $R$, and $K$ is a ring $\operatorname{M}$-ideal of $J$, then it is an exercise that $K$ is a ring $\operatorname{M}$-ideal of $R$ (see Proposition 5.6 for the proof if needed).
Note that a nonunital ring $R$ is an $\operatorname{M}$-ideal of $R^{1}$, since it is an essential ideal. More generally, we have from the observations above that:
Corollary 2.11.
If $R$ is a nonunital ring then the $\operatorname{M}$-ideals of $R$ are the $\operatorname{M}$-ideals of the unitization $R^{1}$ that are contained in $R$. They are also the intersections of $R$ with the $\operatorname{M}$-ideals of $R^{1}$.
Proof.
The first assertion follows from the observations above the corollary. If $J$ is an $\operatorname{M}$-ideal of $R^{1}$ then $J\cap R$ is an $\operatorname{M}$-ideal of $R^{1}$ by Proposition 2.5 (1). Hence, it is an $\operatorname{M}$-ideal of $R$ by the first assertion. ∎
Using this corollary, many of the results in our paper can be extended to $\operatorname{M}$-ideals of nonunital rings in the obvious way. We will not usually take the time to do this, but instead leave it to the reader.
We note that if $J_{k}$ is an $\operatorname{M}$-ideal in $R_{k}$ for $k=1,\ldots,m$, then $\oplus_{k}\,J_{k}$ need not be an $\operatorname{M}$-ideal in $\oplus_{k}\,R_{k}$. Indeed $J_{1}\oplus R_{2}$ need not be an $\operatorname{M}$-ideal in $R_{1}\oplus R_{2}$. Thus, the sum of two $\operatorname{M}$-ideals need not be an $\operatorname{M}$-ideal. Similarly, $J_{1}\otimes J_{2}$ need not be an $\operatorname{M}$-ideal in $R_{1}\otimes R_{2}$, even if $J_{2}=R_{2}$.
Remark 2.12.
Based on the above abundance of examples, one may be tempted to believe that all nonzero ideals of rings are $\operatorname{M}$-ideals. However, this is easily seen not to be true. Consider the simple example of the ring $\mathds{Z}_{30}$ and the ideal $J:=(2)$. If we take $I_{1}:=(3)$ and $I_{2}:=(5)$, then $J\cap I_{1}=(6)$, $J\cap I_{2}=(10)$, and $I_{1}\cap I_{2}=(15)$. However, $J\cap(I_{1}\cap I_{2})=(0)$.
3. Characterizations and Applications
Our next goal is to give a characterization of $\operatorname{M}$-ideals of rings, and this will illuminate further why minimal ideals are $\operatorname{M}$-ideals. This result and its consequences are an ideal-theoretic adaptation from [7]. Before we state the result, we need a definition.
Definition 3.1.
We say that an ideal $I$ of a ring $R$ is relatively irreducible111cf. An ideal $I$ of $R$ is called irreducible if for any two ideals $J$, $K$ of $R$ and $I=J\cap K$ implies that either $J=I$ or $K=I$. Our ‘relatively irreducible’ is a form of ‘prime-ness’ if for every ideal $J$ and $K$ of $R$ with $J\subseteq I$, $K\subseteq I$, and $J\cap K=0$ implies that either $J=0$ or $K=0$.
Theorem 3.1.
An ideal $I$ of a ring $R$ (not necessarily unital) is an $\operatorname{M}$-ideal if and only if either it is essential or relatively irreducible.
Proof.
Clearly, if $I$ is essential, then $I$ is an $\operatorname{M}$-ideal by Proposition 2.3. Now suppose that $I$ is relatively irreducible. Let $J$ and $K$ be two nonzero ideals of $R$ such that $J\cap K\neq 0$, $I\cap J\neq 0$, and $I\cap K\neq 0$. Since $I$ is relatively irreducible, this implies
$I\cap(J\cap K)=(I\cap J)\cap(I\cap K)\neq 0.$
Hence $I$ is an $\operatorname{M}$-ideal. Conversely, suppose that $I$ is an $\operatorname{M}$-ideal and $I$ is not essential. We show that $I$ is relatively irreducible. Since $I$ is not essential, there exists a nonzero ideal $L$ in $R$ such that $I\cap L=0$. Let $J$ and $K$ be nonzero ideals of $R$ such that $J\subseteq I$ and $K\subseteq I$. We claim that $J\cap K\neq 0$. Let $J^{\prime}:=J+L$ and $K^{\prime}:=K+L$. Then $J^{\prime}\cap K^{\prime}\supseteq L\neq 0$. Applying modular law, we obtain
$I\cap J^{\prime}=I\cap(J+L)=J+(I\cap L)=J\neq 0.$
Similarly, we get $I\cap K^{\prime}=K\neq 0$. Since $I$ is a $\operatorname{M}$-ideal, we have
$J\cap K=(I\cap J^{\prime})\cap(I\cap K^{\prime})=I\cap(J^{\prime}\cap K^{\prime})\neq 0.\qed$
Remark 3.2.
Every minimal ideal is relatively irreducible. Indeed, let $I$ be a minimal ideal of $R$. Suppose that $J\subseteq I$, $K\subseteq I$, and $J\cap K=0$. Since $I$ is minimal, the only case we need to care about is when $J=I$ and $K=I$. But then $J\cap K=I\neq 0$, a contradiction. Thus, the assertion in Proposition 2.5(6) is no longer surprising. However, note that there exist $\operatorname{M}$-ideals that are neither essential nor minimal (for example, consider the ideal $3\mathds{Z}_{12}$ in the ring $\mathds{Z}_{12}$). So, Theorem 3.1 can not be improved further.
For the rest of this section, we shall see some applications of the above characterization theorem. Recall that the sum of minimal ideals of a ring $R$ is called the socle of $R$, and we denote it by $\mathrm{Soc}(R)$. It is well-known that the socle of a ring $R$ is the intersection of all essential ideals of $R$. Now, Theorem 3.1, tells us when $\mathrm{Soc}(R)$ is an $\operatorname{M}$-ideal.
Theorem 3.3.
[…]
Proof.
If $L$ has no minimal ideals, then $\mathrm{Soc}(L)$ is trivially an $\operatorname{M}$-ideal. If $R$ has only one minimal ideal $I$, then $\mathrm{Soc}(L)=I$, is relatively irreducible and hence, by Theorem 3.1, is an $\operatorname{M}$-ideal. If $R$ has more than one minimal ideal, then $\mathrm{Soc}(L)$ is not relatively irreducible, as the intersection of any two distinct minimal ideals $I$ and $J$ is 00. Indeed: $I\cap J\subseteq I$, whence $I\cap J=0$ or $I\cap J=I$. But if $I\cap J=I$, then $I\subseteq J$, whence $I=J$. Therefore, if $R$ has more than one minimal ideal, then by Theorem 3.1, $\mathrm{Soc}(R)$ is an $\operatorname{M}$-ideal precisely when it is essential. ∎
Using Theorem 3.1, we shall now characterize $\operatorname{M}$-ideals of rings $\mathds{Z}_{n}$.
Theorem 3.4.
Consider a ring $\mathds{Z}_{n}$ with $n=p_{1}^{m_{1}}\cdot\ldots\cdot p_{k}^{m_{k}}$. Then a nontrivial ideal
$I:=\left(p_{1}^{m^{\prime}_{1}}\cdot\ldots\cdot p_{k}^{m^{\prime}_{k}}\right)$
is an $\operatorname{M}$-ideal if and only if there are not $i$, $j$, $s\in\{1,\ldots,k\}$ such that $i\neq j$, $m^{\prime}_{i}<m_{i}$, $m^{\prime}_{j}<m_{j}$, and $m^{\prime}_{s}=m_{s}$. In particular, $I$ is essential if and only if for every $i$ we have $m^{\prime}_{i}<m_{i}$.
Proof.
We first prove that $I$ is essential if and only if for every $i$ we have $m^{\prime}_{i}<m_{i}$. If $m^{\prime}_{i}<m_{i}$, given an ideal $J\neq 0$ with $J=(p_{1}^{m^{\prime\prime}_{1}}\cdot\ldots\cdot p_{k}^{m^{\prime\prime}_{k}})$, we obtain
$I\cap J=\left(p_{1}^{\operatorname{max}\{m^{\prime}_{1},m^{\prime\prime}_{1}\}}\cdot\ldots\cdot p_{k}^{\operatorname{max}\{m^{\prime}_{k},m^{\prime\prime}_{k}\}}\right).$
Since $J\neq 0$, there exists $j\in\{1,\ldots,k\}$ such that $m^{\prime\prime}_{j}<m_{j}$ and so $I\cap J\neq 0$ since $p_{j}$ appears with exponent $\operatorname{max}\{m^{\prime}_{j},m^{\prime\prime}_{j}\}<m_{j}$ in the unique factorization of the generator of $I\cap J$. Conversely, if $I$ is essential, then for every $i\in\{1,\ldots,k\}$ we must have
$I\cap\left(p_{1}^{m_{1}}\cdot\ldots\cdot p_{i}^{m_{i}-1}\cdot\ldots\cdot p_{k}^{m_{k}}\right)\neq 0,$
and thus $\operatorname{max}\{m_{i}-1,m^{\prime}_{i}\}<m_{i}$. So, it must be $m^{\prime}_{i}<m_{i}$.
Let us now consider the case in which $I$ is not essential, and thus there exists $s\in\{1,\ldots,k\}$ such that $m^{\prime}_{s}=m_{s}$. Suppose there are not $i$, $j\in\{1,\ldots,k\}$ such that $i\neq j$, $m^{\prime}_{i}<m_{i}$, $m^{\prime}_{j}<m_{j}$. If $m^{\prime}_{j}=m_{j}$ for every $j\neq i$ and $m^{\prime}_{i}=m_{i}-1$, then $I$ is a minimal ideal and thus, an $\operatorname{M}$-ideal. Indeed, if an ideal
$J=\left(p_{1}^{m^{\prime\prime}_{1}}\cdot\ldots\cdot p_{k}^{m^{\prime\prime}_{k}}\right)$
[…]
Suppose now that there exists $i,j\in\{1,\ldots,k\}$ such that $i\neq j$, $m^{\prime}_{i}<m_{i}$, $m^{\prime}_{j}<m_{j}$ and consider the following ideals:
$J=\left(p_{1}^{m_{1}}\cdot\ldots\cdot p_{i}^{m_{i}}\cdot\ldots\cdot p_{j}^{m^{\prime}_{j}}\cdot\ldots\cdot p^{0}_{s}\cdot\ldots\cdot p_{k}^{m_{k}}\right)\quad\text{ and }\quad L=\left(p_{1}^{m_{1}}\cdot\ldots\cdot p_{i}^{m^{\prime}_{i}}\cdot\ldots\cdot p_{j}^{m_{j}}\cdot\ldots\cdot p^{0}_{s}\cdot\ldots\cdot p_{k}^{m_{k}}\right).$
Then, since $m_{s}>0$, we have
$J\cap L=\left(p_{1}^{m_{1}}\cdot\ldots\cdot p_{i}^{m_{i}}\cdot\ldots\cdot p_{j}^{m_{j}}\cdot\ldots\cdot p^{0}_{s}\cdot\ldots\cdot p_{k}^{m_{k}}\right)\neq 0.$
Moreover, since since $m^{\prime}_{j}<m_{j}$, we have
$I\cap J=\left(p_{1}^{m_{1}}\cdot\ldots\cdot p_{i}^{m_{i}}\cdot\ldots\cdot p_{j}^{m^{\prime}_{j}}\cdot\ldots\cdot p_{s}^{m_{s}}\cdot\ldots\cdot p_{k}^{m_{k}}\right)\neq 0$
and analogously, $I\cap L\neq 0$ because $m^{\prime}_{i}<m_{i}$. But clearly $I\cap J\cap L=0$. So, $I$ is not an $\operatorname{M}$-ideal. ∎
Remark 3.5.
Let us now apply the above result to show that the product of $\operatorname{M}$-ideals may not be an $\operatorname{M}$-ideal. Consider the ring $\mathds{Z}_{900}$. Since $900=2^{2}\cdot 3^{2}\cdot 5^{2}$, by Theorem 3.4, the ideals $(6)=(2\cdot 3)$ and $(10)=(2\cdot 5)$ are $\operatorname{M}$-ideals. However, again by Theorem 3.4, $(6)\cdot(10)=(60)=(2^{2}\cdot 3\cdot 5)$ is not a $\operatorname{M}$-ideal.
Also, notice that a maximal ideal of a ring may not be an $\operatorname{M}$-ideal. Consider the ring $\mathds{Z}_{180}$. The ideal $(5)$ is maximal but is not an $\operatorname{M}$-ideal by Theorem 3.4. Also, the Jacobson radical, $\mathrm{Jac}(R)$, may not be an $\operatorname{M}$-ideal. For example, $\mathrm{Jac}(\mathds{Z}_{180})=(2\cdot 3\cdot 5)=(30)$, is not a $\operatorname{M}$-ideal, again thanks to Theorem 3.4.
Let $C$ be an $R$-module and $A$ be a submodule of $C$. Recall from [6] that $A$ is said to be essentially closed in $C$ provided $A$ has no proper essential extensions within $C$. Motivated by this, we now introduce a similar notion in our context and show its relation with $\operatorname{M}$-ideals.
Definition 3.2.
We say an ideal $I$ is called essentially closed (resp. $\operatorname{M}$-closed) in a ring $R$, if $I$ is an essential ideal (resp. an $\operatorname{M}$-ideal) in an ideal $J$ of $R$, then $I=J$.
Theorem 3.6.
Let $R$ be a ring and let $I$ be a proper ideal of $R$. The following are equivalent:
  1. (1)
    $I$ is $\operatorname{M}$-closed;
  2. (2)
    $I$ is essentially closed and not relatively irreducible.
Proof.
(1$\Rightarrow$(2): Let $I$ be $\operatorname{M}$-closed. Since every essential ideal is an $\operatorname{M}$-ideal, clearly $I$ is essentially closed. Moreover, $I$ is not irreducible. Indeed, if $I$ is irreducible in $R$, by Theorem 3.1, $I$ is an $\operatorname{M}$-ideal of $R$. Since $I\neq R$, this implies that $I$ is not $\operatorname{M}$-closed, which is a contradiction. Hence, $I$ is not relatively irreducible in $R$.
(2$\Rightarrow$(1): Assume that $I\neq R$ is essentially closed and not relatively irreducible. Let $J$ be an ideal of $R$ such that $I$ is an $\operatorname{M}$-ideal of $J$. We show that $J=I$. By Theorem 3.1, $I$ is either relatively irreducible in $J$ or it is essential in $J$. If $I$ is relatively irreducible in $J$, then $I$ is relatively irreducible in $R$ as well, which is a contradiction. If $I$ is essential in $J$, then $I=J$ because $I$ is essentially closed. Hence, $I$ is $\operatorname{M}$-closed. ∎
We shall now see an example of a ring with both essential and minimal $\operatorname{M}$-ideals.
Theorem 3.7.
The nonzero ring $\operatorname{M}$-ideals in $C(K)$ are precisely either the essential ideals or the minimal ideals singly generated by the characteristic function of an isolated point.
Proof.
As noted in Remark 2.2, the closed ideals, which are the Banach space $\operatorname{M}$-ideals of the ring $C(K)$, are the subspaces $J_{E}$ of functions vanishing on a closed subset $E\subset K$. If $E=K\setminus\{x\}$ for $x\in K$ then $J_{E}$ satisfies the conditions of Proposition 2.5(6), so is an $\operatorname{M}$-ideal. Let $E\subset K$ be a closed subset such that $J_{E}$ is a ring $\operatorname{M}$-ideal. If $E=K\setminus U$ and $U$ is not a singleton (the case above), then we claim that $U$ is dense. For if not, then there is a nonempty open set $W\subset E$. Any points $x\neq y$ in $U$ have disjoint neighborhoods $W_{1},W_{2}$ inside $U$. Set $V_{k}=W\cup W_{k}$, and $E_{k}=K\setminus V_{k}$, with $I_{k}=J_{E_{k}}$. Then $\cup_{k}\,E_{k}\neq K$, so that $\bigcap_{k=1}^{n}I_{k}\neq 0$, and $E\cup E_{k}\neq K$ for each $k$, but $E\cup\cup_{k}\,E_{k}=K$. That is, $I\cap\bigcap_{k=1}^{n}I_{k}=0$. Thus, $U$ is dense.
It is known [8] that a nonzero ideal $I$ in $C(K)$ is essential if and only if $\bar{I}$ is essential and if and only if the interior of $Z(I)=Z(\bar{I})$ is empty. Here
$Z(I)=\bigcap_{f\in I}\,f^{-1}(\{0\}).$
If $\bar{I}=J_{E}$ then $E=Z(\bar{I})$, and the interior of $E$ is empty if and only if $U=K\setminus E$ is dense. See also II.5.4.7 in [3]. If $I$ is a ring $\operatorname{M}$-ideal in $C([0,1])$ which is not of the form in the first paragraph of this proof, then by the above, $\bar{I}$ is an essential ideal, so that $I$ is an essential ideal.
Thus, the nonzero ring $\operatorname{M}$-ideals in $C(K)$ are the essential ideals or the minimal ideals singly generated by the characteristic function of an isolated point. ∎
Thus, for example, the ring $\operatorname{M}$-ideals in $C([0,1])$ are exactly the essential ideals. This is because $[0,1]$ has no isolated points.
4. In C*-algebras
The corresponding problem for a noncommutative unital C*-algebra $B$ is also interesting. As a first step, one can show that an ideal J in $B$ is essential in the algebraic sense if and only if $\bar{J}$ is essential in the C*-sense. Indeed, this follows very easily from:
Lemma 4.1.
For nonzero two-sided ideals $I,J$ in a C*-algebra $B$ we have $\overline{I\cap J}=\bar{I}\cap\bar{J}$.
Proof.
Any contractive approximate identity (cai) for $\bar{I}$ may be approximated to give a cai $(e_{t})$ for $I$ which is also a cai for $\bar{I}$. If $x\in\bar{I}\cap\bar{J}$, with bounded approximation $j_{s}\to x$ with $j_{s}\in J$, then
$x=\lim_{t}e_{t}x=\lim_{t}\lim_{s}\,e_{t}j_{s}.$
This is in $\overline{I\cap J}$, since
$e_{t}j_{s}\in IJ\subseteq I\cap J\subseteq\overline{I\cap J}.$
It follows that $\overline{I\cap J}=\bar{I}\cap\bar{J}$. ∎
The same proof works in a Banach algebra, provided that $\bar{I}$ has a bounded approximate identity. This must be well-known.
We recall that a C*-algebra $A$ is prime if it does not contain two proper closed orthogonal two-sided ideals. For example, $B(l^{2})$ is prime since it has a unique closed two-sided ideal. If $A$ contains two nonzero central positive elements with product 0, then $A$ is not prime. However, this condition does not characterize (non)primeness. Rather, $A$ is prime if and only if for all nonzero $x$, $y\in A$ we have $xAy\neq 0$ [3, Proposition II.5.4.5]. Also, $A$ is prime if and only if every nonzero ideal of $A$ is an essential ideal. In the separable case, primeness is equivalent to the existence of a faithful irreducible representation [3].
Proposition 4.2.
For a nonzero ideal $J$ in a C*-algebra the following conditions are equivalent:
  1. (1)
    $J$ is relatively irreducible.
  2. (2)
    $\bar{J}$ is a prime C*-algebra
  3. (3)
    $xJy\neq 0$ for nonzero $x,y\in J$.
Proof.
For a closed ideal in a C*-algebra, ‘relatively irreducible’ is precisely the same as being a prime C*-algebra (see [3, Definition II.5.4.4]). It follows that for a non-closed ideal $J$ with $\bar{J}$ a prime C*-algebra, Lemma 4.1 implies that $J$ is relatively irreducible. Indeed for nonzero ideals $I,K$ in $J$ with $I\cap K=0$ we have $\bar{I}\cap\bar{K}=0$, giving a contradiction. So (2) implies (1).
The equivalence of (2) and (3) uses [3, Proposition II.5.4.5] and Lemma 4.1. If $xJy=0$, then by continuity $x\bar{J}y=0$, while if $x\bar{J}y=0$ then of course $xJy=0$. Hence, if $\bar{J}$ is a prime C*-algebra, then $xJy\neq 0$ for nonzero $x,y\in J$. Conversely, suppose that (3) holds. Suppose that $I,K$ are nonzero closed ideals in $\bar{J}$. Then $I\cap J\neq 0$ or else Lemma 4.1 gives a contradiction. Similarly, $K\cap J\neq 0$. Suppose that $0\neq x\in I\cap J,0\neq y\in K\cap J$. Then $x\bar{J}y\subset I\cap K$. If $I\cap K=0$ we obtain the contradiction that $x$ or $y$ is 0. Thus $I\cap K\neq 0$. Thus $\bar{J}$ is a prime C*-algebra.
Finally, suppose that (1) holds. The argument above shows that if $I,K$ are nonzero closed ideals in $\bar{J}$ then $I\cap J\neq 0$ and $K\cap J\neq 0$. So $I\cap K\cap J=0$. By Lemma 4.1 we have $I\cap K\cap\bar{J}=0$. So $\bar{J}$ is a prime C*-algebra, and we have (2). ∎
The following result now follows from Theorem 3.1.
Corollary 4.3.
The nonzero ring $\operatorname{M}$-ideals in a C*-algebra $B$ are precisely either the essential ideals or the ideals $J$ whose closure is a prime C*-algebra. Such ideals $J$ are exactly the ideals such that $xJy\neq 0$ for nonzero $x$, $y\in J$.
Corollary 4.4.
If $I$ is a ring $\operatorname{M}$-ideal in a unital C*-algebra $A$ and if $I\subseteq B\subseteq A$ with $B$ a subalgebra, then $I$ is a ring $\operatorname{M}$-ideal in $B$.
Proof.
If $\bar{I}$ is a prime C*-algebra then $I$ is relatively irreducible by the proposition, and hence is a ring $\operatorname{M}$-ideal in $B$. On the other hand, if $I$ is essential in $A$, then so is $\bar{I}$. The left regular representation on $\bar{I}$ is faithful on $A$ (this is a well known characterization of essential ideals in C*-algebras, and is a nice exercise in that subject). Hence it is also faithful on the closure $\bar{B}$ of $B$. However this easily implies that $\bar{I}$ is essential in $\bar{B}$.
The argument for Lemma 4.1 now works to show that for a nonzero ideal $J$ in $B$ we have $\overline{I\cap J}=\bar{I}\cap\bar{J}\neq 0$ in $\bar{B}$, and so $I\cap J\neq 0$. That is, $I$ is essential in $B$. ∎
We conclude this section with a couple of examples of $\operatorname{M}$-ideals in some famous nonselfadjoint operator algebras and function algebras.
Proposition 4.5.
Every ideal in the upper triangular matrices $\mathcal{T}_{n}$, or in in the disk algebra or in $H^{\infty}(\mathds{D})$, is an $\operatorname{M}$-ideal.
Proof.
Every ideal in the upper triangular matrices $\mathcal{T}_{n}$ is essential and hence is an $\operatorname{M}$-ideal. This is because every nonzero ideal $J$ in $\mathcal{T}_{n}$ contains $e_{1n}$ (this is easily seen by looking at $e_{1i}Je_{jn}$).
Similarly, every nonzero ideal in the disk algebra or in $H^{\infty}(\mathds{D})$ is essential and hence is an $\operatorname{M}$-ideal. Indeed, this follows from Proposition 2.5 since these are integral domains. If $I$ and $J$ are nonzero ideals here, then their intersection is nonzero. This follows since for nonzero $f\in I$ and $g\in J$ we must have $fg$ not identically 00 on the disk. ∎
5. Some Structural Aspects
We shall now focus on studying the influence of $\operatorname{M}$-ideals on the structure of rings. We recall that a ring $R$ is subdirectly irreducible if it is not expressible as a subdirect product of two nonzero rings (see, e.g., [10]). This is equivalent to all the non-zero ideals having a non-zero intersection. It is also equivalent to the existence of an element which is a multiple of all non-zero elements of the ring.
Proposition 5.1.
Suppose that a unital ring $R$ has an ideal decomposition $R=I\oplus J$. Then $I$ is an $\operatorname{M}$-ideal in $R$ if and only if every finite collection of nonzero ideals $I_{1},I_{2},\dots,I_{n}$ of $R$ has $\bigcap_{k=1}^{n}I_{k}\neq 0$. (This includes the case that $I$ is simple, known as the ‘vacuous case’.)
Proof.
If $I$ has $n$ ideals $J_{k}$ with intersection $(0)$ set $I_{k}=J\oplus J_{k}$. We have $\cap_{k}I_{k}=J$. Indeed, suppose that $x\in\bigcap_{k}I_{k}$ with $x=y_{k}+j_{k}$ for all $k$ and $y_{k}\in J,j_{k}\in J_{k}$. Then all of the $y_{k}$ are the same, so that $y_{k}=0$ by the assumption at the start of the proof. So $x\in J$. We deduce that $I\cap(\bigcap_{k}I_{k})=0$, so that $I$ is not an $\operatorname{M}$-ideal.
Conversely, suppose that $I$ is not an $\operatorname{M}$-ideal. Then there exist nonzero ideals $I_{1}$ and $I_{2}$ of $R$ with $\bigcap_{k=1}^{2}I_{k}\neq 0$ and $I\cap I_{k}\neq 0$ for each $k$, but $\bigcap_{k}\,(I\cap I_{k})=I\cap(\bigcap_{k}I_{k})=0$. Thus, we have found a finite collection of nonzero ideals of $I$ with zero intersection. ∎
Our next theorem characterizes direct sum simple-representation of rings in terms of the absence of $\operatorname{M}$-ideals, and it extends Theorem 3 from [11]. To prove the result, we first need a lemma.
Lemma 5.2.
If $I$ is a nonzero proper ideal of a ring $R$, then either $I$ is a direct summand of $R$ (so is unital and has a complementary ideal in $R$), or there exists a proper $\operatorname{M}$-ideal of $R$ containing $I$.
Proof.
Suppose that $I$ is an ideal of $R$ which is not a direct summand. We claim that there exists an $\operatorname{M}$-ideal $K$ of $R$ such that $I\subseteq K$. Let $\mathcal{C}$ be the collection of all ideals $K$ of $R$, which are not a direct summand of $R$, such that $I\subseteq K$ and $K\neq R$. Since $I\in\mathcal{C}$, the set $\mathcal{C}$ is nonempty. By Zorn’s Lemma, the set $\mathcal{C}$ is partially ordered by inclusion, and every chain in $\mathcal{C}$ has an upper bound. Note that this upper bound, the union of the chain, cannot be a direct summand of $R$. For if it were, then it and hence one of the ideals $J_{i}$ in the chain, is unital. Hence, $J_{i}$ is a direct summand, which is impossible. Therefore, $\mathcal{C}$ has a maximal element $K$.
To prove that $K$ is an $\operatorname{M}$-ideal, let $I_{1}$ and $I_{2}$ be nonzero ideals of $R$ such that $I_{1}\cap I_{2}\neq 0$ and $K\cap I_{k}\neq 0$ for each $k$. If $K\cap\big{(}\bigcap_{k=1}^{2}I_{k}\big{)}=0$, then the ideal $K+\big{(}\bigcap_{k=1}^{2}I_{k}\big{)}$ is strictly larger than $K$ containing $I$, and
$K+\big{(}\bigcap_{k=1}^{2}I_{k}\big{)}\neq R,$
since $K$ is not a direct summand of $R$. A similar argument shows that $K+\big{(}\bigcap_{k=1}^{2}I_{k}\big{)}$ is not a direct summand. This contradicts the maximality of $K$ in $\mathcal{C}$. Hence, $K\cap\big{(}\bigcap_{k=1}^{2}I_{k}\big{)}\neq 0$, proving that $K$ is an $\operatorname{M}$-ideal. ∎
For example, suppose that not every ideal in $I$ is essential. So $I$ has two ideals $J_{1}$, $J_{2}$ with intersection 00. Then $I$ is not an $\operatorname{M}$-ideal of $R$ by the last result.
Theorem 5.3.
For any ring $R$, the following are equivalent:
  1. (1)
    Every proper $\operatorname{M}$-ideal of $R$ is a direct summand of $R$.
  2. (2)
    Every proper $\operatorname{M}$-ideal of $R$ is a simple ideal of $R$ which is also a direct summand.
  3. (3)
    Every proper ideal of $R$ is a direct summand of $R$.
  4. (4)
    $R$ is a direct sum of simple rings.
Proof.
If every $\operatorname{M}$-ideal of $R$ is a direct summand of $R$, then $R$ has no essential ideals. By Proposition 2.3 and Lemma 1 from [11], every ideal is a direct summand of $R$. Hence, (1) implies (3). Clearly (2) implies (1). For a proof of (3$\Leftrightarrow$(4), we refer to [1].
If (4) holds and $J$ is a proper ideal in $R$, then it is easy to see that $J$ is the direct sum $\oplus_{i\in I}J_{i}$ of certain of the simple subrings. Let $J_{0}$ be one of the simple subrings different from the $J_{i}$. Suppose that $I$ is not singleton, with $i,j\in I$, $i\neq j$. Setting $I_{1}=J_{0}\cup J_{i},I_{2}=J_{0}\cup J_{j}$, we have $\bigcap_{k=1}^{n}I_{k}\neq 0$, and $J\cap I_{k}\neq 0$ for each $k$, but
$J\cap\left(\bigcap_{k=1}^{n}I_{k}\right)=0.$
So, $J$ is not an $\operatorname{M}$-ideal. On the other hand, we know simple ideals are $\operatorname{M}$-ideals. This proves (2). ∎
The conditions above are not equivalent to: every proper $\operatorname{M}$-ideal of $R$ is a simple ideal of $R$. For example, let $R$ be the unitization of $M_{n}$. This has the latter property (it has one proper $\operatorname{M}$-ideal, but there are no proper ideals that have a complemented ideal).
Corollary 5.4.
A ring with no proper $\operatorname{M}$-ideals is simple.
Proof.
By the lemma, every proper ideal is a direct summand. So by the theorem, $R$ is a direct sum of simple rings. Since simple ideals are $\operatorname{M}$-ideals, $R$ must in fact be simple. ∎
Theorem 5.5.
Morita equivalence of unital algebras preserves essential ideals and $\operatorname{M}$-ideals.
Proof.
Suppose that $(R,S,X,Y)$ is a Morita context with $X\otimes_{R}Y\cong S$ and $Y\otimes_{S}X\cong R$. In particular, we may write $1_{R}=\sum_{k=1}^{n}(y_{k},x_{k})$ and $1_{S}=\sum_{k=1}^{m}[w_{k},z_{k}]$, with, $x_{k},w_{k}\in X,$ and $y_{k},z_{k}\in Y$. The map from ideals $J$ in $S$ to ideals in $R$, for example, takes $\{a\in J\}$ to
$T(J)=\left\{\sum_{k=1}^{N}(r_{k},as_{k})\mid a\in J,N\in\operatorname{{\mathds{N}}},r_{k}\in Y,s_{k}\in X\right\}.$
It is well known that $T(J\cap K)=T(J)\cap T(K)$; indeed, Morita equivalence induces a lattice isomorphism of the ideal lattice of $S$ onto the ideal lattice of $R$. Moreover, $J\cap K=(0)$ if and only if $T(J)\cap T(K)=(0)$. Thus, $J$ is essential if and only if $T(J)$ is essential. Similarly, $J$ is an $\operatorname{M}$-ideal if and only if $T(J)$ is an $\operatorname{M}$-ideal. ∎
The following result is well-known (see, e.g., Proposition 5.6 in [6]) to hold for essential submodules.
Proposition 5.6.
Suppose that $R$ is a ring (not necessarily unital).
  1. (1)
    If $N$ and $N^{\prime}$ are ideals of $R$ such that $N$ is an $\operatorname{M}$-ideal of $N^{\prime}$ and $N^{\prime}$ is an $\operatorname{M}$-ideal of $R$, then $N$ is an $\operatorname{M}$-ideal of $R$.
  2. (2)
    Suppose that $N$ and $N^{\prime}$ are ideals of $R$ such that $N$ is an ideal of $N^{\prime}$. Then if $N$ is an $\operatorname{M}$-ideal of $R$, then $N$ is an $\operatorname{M}$-ideal in $N^{\prime}$.
  3. (3)
    Let $A_{1},A_{2},B_{1},B_{2}$ be ideals of $R$. If $A_{1}$ is an $\operatorname{M}$-ideal of $B_{1}$ and $A_{2}$ is an $\operatorname{M}$-ideal of $B_{2}$, then $A_{1}\cap A_{2}$ is an $\operatorname{M}$-ideal of $B_{1}\cap B_{2}$.
Proof.
(1) Suppose that $N_{1}$ and $N_{2}$ are nonzero ideals of $R$ such that $N_{1}\cap N_{2}\neq 0.$ We assume that $N\cap N_{k}\neq 0$ for each $k$. Then $N^{\prime}\cap N_{k}\neq 0$ for each $k$. Since $N^{\prime}$ is an $\operatorname{M}$-ideal of $R$, we have $N^{\prime}\cap\left(N_{1}\cap N_{2}\right)\neq 0$. This implies that $0\neq N^{\prime}\cap N_{k}\subseteq N^{\prime}$ for each $k$ with $\bigcap_{k=1}^{2}(N^{\prime}\cap N_{k})\neq 0.$ Since $N$ is an $\operatorname{M}$-ideal of $N^{\prime}$, moreover, we have
$0\neq N\cap\left(\bigcap_{k=1}^{2}(N^{\prime}\cap N_{k})\right)=N\cap\left(N_{1}\cap N_{2}\right),$
which proves the claim.
(2) We leave this as an exercise.
[…]
$\bigcap_{k=1}^{2}N_{k}=\bigcap_{k=1}^{2}(N_{k}^{1}\cap N_{k}^{2}).$
Since $(A_{1}\cap A_{2})\cap N_{k}\neq 0$ for each $k$, we get
$0\neq(A_{1}\cap A_{2})\cap N_{k}=(A_{1}\cap A_{2})\cap(N_{k}^{1}\cap N_{k}^{2})=(A_{1}\cap N_{k}^{1})\cap(A_{2}\cap N_{k}^{2}).$
This implies that $A_{1}\cap N_{k}^{1}\neq 0$ and $A_{2}\cap N_{k}^{2}\neq 0$. Since $A_{1}$ is an $\operatorname{M}$-ideal of $B_{1}$ and $A_{2}$ is an $\operatorname{M}$-ideal of $B_{2}$, we have
$A_{1}\cap\left(\bigcap_{k=1}^{2}N_{k}^{1}\right)\neq 0\quad\text{and}\quad A_{2}\cap\left(\bigcap_{k=1}^{2}N_{k}^{2}\right)\neq 0.$
Therefore,
$(A_{1}\cap A_{2})\cap\left(\bigcap_{k=1}^{2}N_{k}\right)=(A_{1}\cap A_{2})\cap\left(\bigcap_{k=1}^{2}(N_{k}^{1}\cap N_{k}^{2})\right)\neq 0.\qed$
Motivated by essential complements of submodules, we shall now introduce a similar notion for $\operatorname{M}$-ideals and study some of their properties.
Definition 5.1.
An $\operatorname{M}$-complement of an ideal $N$ of a ring $R$ is an ideal $N^{\prime}$ of $R$ such that $N\cap N^{\prime}=0$ and $N+N^{\prime}$ is an $\operatorname{M}$-ideal of $R$.
Remark 5.7.
Note that $\operatorname{M}$-complements are not unique. Indeed, in $\mathds{Z}_{12}$, the ideals $(3)$ and $(4)$ are complements of each other, as $(3)\cap(4)=(0)$ and $(3)+(4)=\mathds{Z}_{12}$. Consider $(2)$, $(3)\cap(2)=(6)$, and $(4)\cap(2)=(4)$. It is easily seen that $(2)$ is an $\operatorname{M}$-ideal. Moreover, $(3)\cap(4)=(0)$ and $(6)+(4)=(2)\mathds{Z}_{12}$. Therefore, both $(3)$ and $(6)$ are $\operatorname{M}$-complements of $(4)$.
Proposition 5.8.
If $A$ and $N$ are ideals of an arithmetic ring $R$ such that $A\subseteq N$, and $B$ is an $\operatorname{M}$-complement of $A$ in $R$, then $B\cap N$ is an $\operatorname{M}$-complement of $A$ in $N$.
Proof.
Since $A\cap B=0$ and $B\cap N\subseteq B$, it follows that $A\cap(B\cap N)\subseteq A\cap B=0.$ To show $A+(B\cap N)$ is an $\operatorname{M}$-ideal of $N$, let $K_{1}$ and $K_{2}$ be nonzero ideals of $N$ such that $K_{1}\cap K_{2}\neq 0$ and $(A+(B\cap N))\cap K_{i}\neq 0$, for each $i$. Since $A+(B\cap N)\subseteq A+B$ and $K_{i}\subseteq N\subseteq R$, by applying distributivity twice, we have
$0\neq(A+(B\cap N))\cap K_{i}=(A+B)\cap K_{i},$
for each $i$. Thus, we can rewrite
$(A+(B\cap N))\cap\left(K_{1}\cap K_{2}\right)=(A+B)\cap\left(K_{1}\cap K_{2}\right).$
Since $K_{1}\cap K_{2}\neq 0$ and $(A+B)\cap K_{i}\neq 0$ for each $i$, and $A+B$ is an $\operatorname{M}$-ideal, we obtain $(A+B)\cap\left(K_{1}\cap K_{2}\right)\neq 0.$ Therefore,
$(A+(B\cap N))\cap\left(K_{1}\cap K_{2}\right)\neq 0.\qed$
Theorem 5.9.
If $N$ and $Q$ are ideals of a ring $R$ with $N\cap Q=0$, then $N$ has an $\operatorname{M}$-complement containing $Q$.
Proof.
Consider the set
$\mathcal{S}:=\left\{N^{\prime}\mid N^{\prime}\;\text{is an ideal of}\;R,\,N\cap N^{\prime}=0,\;\text{and}\;Q\subseteq N^{\prime}\right\}.$
Clearly $Q\in\mathcal{S}$, so $\mathcal{S}$ is nonempty. By Zorn’s Lemma, $\mathcal{S}$ has a maximal element, say $N^{\prime}$. We claim that $N+N^{\prime}$ is an $\operatorname{M}$-ideal of $R$. If not, then there exist nonzero ideals $N_{1}$ and $N_{2}$ of $R$ such that $N_{1}\cap N_{2}\neq 0$ and $(N+N^{\prime})\cap N_{k}\neq 0$ for each $k$, but
$(N+N^{\prime})\cap\left(N_{1}\cap N_{2}\right)=0.$
Then by Lemma 6.3 from [5], we have $N\cap\left(N^{\prime}+(N_{1}\cap N_{2})\right)=0,$ contradicted the maximality of $N^{\prime}$. ∎
The following corollary is worth comparing with Theorem 2.4.8 in [12].
Corollary 5.10.
In a ring $R$, every ideal $N$ of $R$ has an $\operatorname{M}$-complement.
Proof.
Take $Q=0$ in Theorem 5.9. ∎
Remark 5.11.
We can also see from Theorem 5.9 that if a ring $R$ does not have any proper $\operatorname{M}$-ideal, then $R$ is complemented. (Indeed: if $R$ does not have any proper $\operatorname{M}$-ideals, then by Corollary 5.10, for every ideal $N$ of $R$, there exists an ideal $N^{\prime}$ of $R$ such that $R=N\oplus N^{\prime}$, and hence, $R$ is complemented.) However, this follows from the stronger result, Corollary 5.4.
6. Further Work
The theory of $\operatorname{M}$-ideals, as developed here, constitutes a generalization of essential ideals and provides a new perspective on structural properties of rings. Their dichotomic characterization of $\operatorname{M}$-ideals—being either essential or relatively irreducible—suggests several avenues for continued investigation, both within classical ring theory and in its broader generalizations.
A natural direction lies in extending the notion of $\operatorname{M}$-ideals to nonassociative algebraic structures, such as Jordan algebras, Lie algebras, or alternative algebras. In these contexts, it would be intriguing to define $\operatorname{M}$-ideals via suitable intersection-theoretic axioms and determine whether an analogous dichotomy persists. This would unify ideal-theoretic notions across a wider class of algebraic systems and may lead to structural results, particularly for algebras with a nontrivial multiplication that is not associative.
Another path concerns the translation of $\operatorname{M}$-ideals into the context of module theory. Just as essential submodules arise naturally in the study of injective hulls, one might define $\operatorname{M}$-submodules via a similar intersection condition among submodules. This opens up the possibility of exploring the existence of $\operatorname{M}$-complements, analogues of essential closures, and their behavior under standard module-theoretic constructions. Questions regarding categorical analogues of M-ideals in abelian categories or Grothendieck categories would further deepen the theory.
Given the important connection between ‘essential’ and ‘injective’ and ‘injective hull’ in both ring theory and functional analysis (see e.g. [4, Chapter 4]), and in particular for Banach and operator space $M$-ideals (see e.g. Section 4.8 of [4] and [9]) it is important to explore this connection both for rings and modules.
Given the topological motivations behind $\operatorname{M}$-ideals in Banach spaces, it is reasonable to investigate their analogues in topological or linearly topological rings. In particular, one may study the behavior of $\operatorname{M}$-ideals under completions, localizations, or limits. It would be worthwhile to determine whether $\operatorname{M}$-ideals are preserved under flat or faithfully flat base change and how they interact with completions in the topological sense, especially in rings of continuous functions or formal power series. Similarly, it would be of interest to study the analogue of our ring $M$-ideals, but where all the ideals involved are taken to be topologically closed. Studying such ideals in common function algebras will often involve peak set arguments.
Further connections may be drawn with radical theory. While Jacobson radicals need not be $\operatorname{M}$-ideals, a systematic study of the interaction between $\operatorname{M}$-ideals and various radicals—such as the upper nilradical, Brown–McCoy radical, or Levitzki radical—may reveal new invariant-theoretic properties. One may define the $\operatorname{M}$-radical of a ring to be the sum or intersection of all its $\operatorname{M}$-ideals, and explore whether it satisfies standard radical-theoretic axioms.
There are further structural aspects related to base change and homomorphisms. For instance, it is already known that $\operatorname{M}$-ideals are preserved under Morita equivalence, yet it remains to investigate their behavior under less restrictive contexts such as tensor extensions, faithfully flat morphisms, or Frobenius maps. The functoriality of $\operatorname{M}$-ideals in diagrammatic or categorical frameworks could also be examined.
A classification of $\operatorname{M}$-ideals in more general classes of rings remains largely open. While the article classifies $\operatorname{M}$-ideals in $\mathds{Z}_{n}$, $C(K)$, and C*-algebras, similar analysis for Dedekind domains, valuation rings, Noetherian rings, PI-rings, or Leavitt path algebras could yield illuminating structural theorems. In many of these settings, one may expect a characterization in terms of central idempotents or minimal prime ideals.
Lastly, it may be fruitful to define invariants associated to $\operatorname{M}$-ideals, such as an “$\operatorname{M}$-dimension” of a ring, defined by the maximal length of a strictly ascending chain of $\operatorname{M}$-ideals, or the “$\operatorname{M}$-length” of a module, measuring the number of nontrivial $\operatorname{M}$-submodules in a composition series. These may provide new quantitative tools for measuring how far a ring is from being $\operatorname{M}$-simple.
Acknowledgements
D.P. Blecher acknowledges support from NSF Grant DMS-2154903, and also from the School of Mathematical and Statistical Sciences at NWU Potchefstroom, which helped make an international visit possible. We also thank Francois Schulz for introducing us to each other.